Phantomdex:

An ice cream shop tracks the orders for different flavors and styles of ice cream to stock the best inventory. The two-way frequency table summarizes the orders for the last week. Hand–dipped Soft Serve Sundae Milkshake Chocolate 25 10 27 33 Vanilla 20 15 43 17 Strawberry 18 12 10 5 Mint Chocolate 12 8 5 10 Determine the empirical probability of a customer ordering a milkshake. 0.650 0.352 0.330 0.241

Phantomdex:

Not Eiwoh scaring me by stalking my question

toga:

To determine the empirical probability of a customer ordering a milkshake, we need to look at the total number of milkshake orders and divide that by the total number of orders. From the given table, we can see that there were 27 + 43 + 10 + 5 = 85 total milkshake orders. And the total number of orders is the sum of all the entries in the table, which is 25 + 10 + 27 + 33 + 20 + 15 + 43 + 17 + 18 + 12 + 10 + 5 + 12 + 8 + 5 + 10 = 260. So the empirical probability of a customer ordering a milkshake is

Phantomdex:

...0.241?

toga:

@phantomdex wrote:
...0.241?
no try again you got this

Phantomdex:

Oh 0.352!

toga:

@phantomdex wrote:
Oh 0.352!
no, it is 0.330

Phantomdex:

Huh? I don't get it-

toga:

@phantomdex wrote:
Huh? I don't get it-
To determine the empirical probability of a customer ordering a milkshake, we need to add up the number of milkshakes ordered and divide by the total number of orders. According to the table, the total number of milkshakes ordered is 27 + 43 + 10 + 5 = 85. The total number of orders is the sum of all the entries in the table, which is 25 + 10 + 27 + 33 + 20 + 15 + 43 + 17 + 18 + 12 + 10 + 5 + 12 + 8 + 5 + 10 = 260. Therefore, the empirical probability of a customer ordering a milkshake is 85/260 = 0.327 or approximately 33%. Therefore, the answer is 0.330.

Phantomdex:

Oh...I was supposed to divide

Phantomdex:

My mistake.

toga:

@phantomdex wrote:
My mistake.
that means your human